ZERO ELECTRON KINETIC ENERGY SPECTRA OF GaNH_{3} AND GaNH_{2}(CH_{3})

\underline{SHENGGANG LI}, GRETCHEN K. ROTHSCHOPF, DINESH PILLAI, BEN WILSON, BRAD SOHNLEIN AND DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Zero electron kinetic energy (ZEKE) spectra were obtained for the GaNH_{3} and GaNH_{2}(CH_{3}) complexes. Ab initio and Franck-Condon factor calculations were performed to help spectral assignments. GaNH_{3} has a ^{1}A_{1} (C_{3v}) ground state in the ion and a ^{2}A’(C_{s}) ground state in the neutral. The ionization potential of the ^{2}A’(C_{s}) state is 40136 (5) cm^{−1}. Metalligand stretching frequencies are 270 cm^{−1} in the ^{1}A_{1} state and 161 cm^{−1} in the ^{2}A’ state. The spectrum of GaNH_{2}(CH_{3}) displays more vibronic transitions but at lower energies than GaNH_{3}. The ionization potential of GaNH_{2}(CH_{3}) is 39331(5) cm^{−1}. The Ga^{+}-N stretching in this case has a frequency of 299 cm^{−1} in the ion. The Ga^{+}-N-C and Ga-N-C bending vibrations have frequencies of 126 and 94 cm^{−1}, respectively.